# Timing and Memory Telemetry on GPUs for AI Governance

### v1

@@ 1. Introduction
Embedding telemetry within Trusted Execution Environments (TEEs) initially appears attractive, as enclaves could in principle protect measurement logic from a compromised host. However, TEEs such as Intel SGX and AMD SEV have been repeatedly broken by practical microarchitectural attacks (e.g., Foreshadow (Van Bulck et al., 2018), CacheOut (Van Schaik et al., 2021)) and fault-based exploits (e.g., Plundervolt (Murdock et al., 2020), SEV-Voltage Glitch (Buhren et al., 2021)). These attacks highlight a persistent mismatch between TEE threat models and real-world adversaries. Similarly, NVIDIA’s Confidential Computing (e.g., H100 CC mode (Corporation, 2023)) does not fully address well-resourced attackers. Recent work has reverse-engineered CC internals (Gu et al., 2025) and demonstrated practical breakages (Chuang et al., 2025) at low cost. Advanced physical and side-channel attacks remain explicitly out of scope (Corporation, 2023), limiting the robustness of CC-based solutions.
This lack of reliable visibility into utilization poses challenges for AI governance and safety. In the absence of trusted hardware telemetry, powerful GPUs can be silently repurposed to train or execute AI models for malicious purposes, including coordinated cyberattacks, automated vulnerability exploitation (Anthropic Team, 2025; Singh, 2025), or advanced malware generation (Metta et al., 2024). At the same time, utilization monitoring must preserve user privacy and avoid leaking sensitive model architectures, training data, or application logic.
@@ 2. GPU in AI Workloads
This motivates telemetry-based utilization measurement, not as a substitute for attestation, but as a complementary signal for enforcement, anomaly detection, or fairness auditing. These mechanisms do not provide cryptographic guarantees; instead, they expose observables such as compute rate, memory bandwidth, or solve-time distributions. Our measurements rely on no hardware trust anchors, but expose timing and memory patterns that could inform future governance protocols.
@@ 3. Untrusted GPU Execution Settings
We consider a range of deployment cases in which external entities may wish to observe or estimate GPU usage for purposes such as telemetry logging, policy enforcement, or future integration into accountability protocols. The goal is to examine whether puzzle-based measurements (e.g., solve times, memory access penalties) provide a useful signal under minimal trust assumptions. We make no claim of full adversarial resistance; the measurements we study are intended as heuristic indicators of usage, not as proofs. To this end, we define four representative settings that differ in which parts of the software/hardware stack can be assumed trustworthy.
@@ 4. Timing Measurements
This section introduces workload-based timing measurements that function as telemetry probes, providing soft-enforcement signals in untrusted environments without TPMs, TEEs, or trusted telemetry. These mechanisms aim to detect whether a worker (e.g., a GPU) has executed a specified unit of computation, when it occurred, and under what degree of contention. These measurements enable latency-based inference, which may inform governance-relevant mechanisms, such as job-scheduling integrity or co-location detection. Each workload probe reveals a different dimension of device behavior: saturation (GEMM), parallel effort (PoW), or sequential pressure (VDF). A randomized challenge schedule enables these probes to be sampled over time, providing the challenger with temporal visibility into resource engagement.
@@ 4.1. Latency Model
Our approach leverages a measurement scheme rooted in memory-hard hash functions. We model PoW solve times under both single-thread and multi-thread (GPU) settings, showing the exponential-time behavior and memoryless property that enables a probabilistic mechanism (see Appendix 8 for more details). This method enforces a relationship between GPU utilization and the computational resources consumed during execution. By employing memory-hard primitives such as Argon2(Biryukov et al., 2016), the scheme ensures that a significant portion of the high-bandwidth memory (HBM) is actively engaged during the measurement process. A practical measurement procedure follows a pattern similar to use cases in cryptocurrencies: the challenger samples a strong random salt from a TRNG, composes a puzzle input that includes the salt and a trusted timestamp, and transmits it to the worker. The worker must perform a large-scale, parallel search or time-bound computation over the input (using thousands of GPU threads) and return a solution within the prescribed time window. The challenger validates the returned solution for correctness and checks that the observed response time lies within the expected interval. Deviations indicate possible under-utilization, outsourcing, or misbehavior. Although this mechanism introduces considerable power and thermal overhead, it provides a quantifiable effort metric of sustained memory utilization within the target GPU, making it well-suited for scenarios with the need for checking resource allocation. Shifting from PoW to GEMM can reduce power consumption and thermal overhead by aligning computation with native tensor-core execution, which is more efficient.
@@
The challenger samples two pseudorandom matrices derived from the salt and timestamp and instructs the worker to perform a sequence of dimension-controlled GEMM operations whose total FLOP count is tightly coupled to the requested utilization level. Because GEMM saturates both compute units and high-bandwidth memory, the resulting latency is highly reproducible on a local GPU but difficult to outsource to CPUs, small GPUs, or remote clouds. Moreover, the challenger can cheaply validate correctness by hashing the returned matrix or verifying random linear combinations of rows and columns (an example of a hash-threshold GEMM puzzle is described in Appendix 10). This GEMM-based PoW thus provides a practical, hardware-aligned mechanism for continuous utilization measurement that exploits the deterministic throughput characteristics of modern GPU matrix engines.
Sequential Delay. Shifting from PoW to Verifiable Delay Function (VDF) (Boneh et al., 2018; Wesolowski, 2020; Rivest et al., 1996) enables the challenger to capture not only parallel effort but also sequential compute pressure, offering finer-grained insight into whether the worker sustained uninterrupted execution over time. This capability is essential in governance scenarios where precomputation or parallel outsourcing must be disincentivized. To this end, we deploy a multi-instance VDF strategy that measures the effective sequential compute capacity of GPUs. Our design is optimized for GPU execution, minimizing kernel divergence and enhancing predictability. In particular, we adopt a construction that balances simplicity and efficiency (Wesolowski, 2020), which is well-suited for continuous utilization monitoring in a multi-instance setting across thousands of GPU cores. Each VDF instance is solved independently and sequentially, ensuring that parallelism provides throughput gains only within controlled architectural limits. At regular intervals, the challenger generates randomized VDF puzzles and transmits each input, including a securely generated salt and timestamp, to the worker. The GPU then performs multi-instance VDF evaluations locally by executing sequential squaring operations over large modular arithmetic groups. Upon completion, the worker returns both the computed outputs and associated proof elements to the challenger; see more theoretical consideration in Appendix 9. The challenger then checks both the correctness of the computation and the timing constraints associated with the response. To mitigate the risk of outsourcing to alternate platforms such as CPU clusters or ASICs, one tunes instance difficulty, modulus size, and iteration depth to favor GPU architectures.
@@ 4.2. Continuous Compute-based Measurement
The challenger continuously sends the worker randomized, time-bound puzzles, enforces an expected utilization level by adjusting difficulty probabilistically, and resists outsourcing by building puzzles that exploit GPU architectural advantages (massive parallelism, tensor cores, and high memory bandwidth) via memory-hard functions, verifiable GEMMs, and multi-instance VDFs. Verification checks both solution validity and timing to detect overuse or policy violations.
@@ 4.3. Random-Timed Challenge-Response
In contrast to the continuous approach, which incurs inherent unavoidable power and computation overhead, we propose a random-timed Challenge-Response mechanism. For this purpose, we focus on the architectural features that are particularassociated with GPUs. Specifically, we build on our approach, leveraging the GPU’s high bandwidth and large memory. Furthermore, we emphasize the heavy memory transitions in today’s LLMs on cutting-edge GPUs and select the memory utilization for measurement.
Figure 2 shows a high-level flow of the proposed challenge–response mechanism, which explicitly exploits the properties of the GPU’s high-bandwidth memory (HBM). In this scheme, a pre-determined challenge data block is stored inside the GPU’s VRAM before the measurement process begins. This part, highlighted in red in Figure 2 and referred to as the Init. Mem., serves as the trusted memory base for all subsequent computations. Any violation of this storage requirement (e.g., moving challenge data off HBM) invalidates the timing assumptions. Our detection method relies on the observation that the computation delay of a given hash function operating on data locally stored in HBM is statistically lower than when the same data must traverse off-chip over PCIe. This latency gap is used to identify compliant versus non-compliant behavior.
@@ 4.4. Comparison of the proposed methods
The CRP method leverages the architectural properties of the GPU’s High Bandwidth Memory (HBM) rather than relying solely on cryptographic assumptions. By verifying whether challenge data resides locally in HBM, it detects off-chip data accesses with minimal delay, achieving low detection time and negligible continuous power overhead. While it primarily offers architectural rather than compute-based enforcement, it provides strong resistance to outsourcing and enables near-instant anomaly detection via timing deviations between on-chip and PCIe-bound operations.
@@ 6.4. Overhead Analysis
This follows directly from the protocol design where the CHAL region must occupy a large, contiguous portion of the GPU’s HBM so that residency violations can be detected with high sensitivity. Consequently, the memory footprint of the VRAM Residency protocol is large by design, even though it performs no additional allocations beyond the CHAL itself.
@@ 7. Conclusion and Discussion
While our fingerprinting-based masking mechanism was evaluated in a non-continuous setting, its integration with randomized probes offers a promising direction for future continuous telemetry mechanisms. Periodic sampling with ephemeral nonces and small Argon2 probes could enforce ongoing residency checks with low overhead, subject to constraints imposed by co-residency, scheduling, and bandwidth fairness. Designing such extensions without disrupting legitimate workloads remains an open challenge. While Section 6 shows separable timing distributions, we do not define formal thresholds or statistical tests. As such, FP/FN rates are not quantified, though they can be estimated empirically under deployment-specific assumptions. A detection-theoretic framework remains important future work.

### v2

@@ 1. Introduction
Embedding telemetry within Trusted Execution Environments (TEEs) initially appears attractive, as enclaves could in principle protect measurement logic from a compromised host. However, TEEs such as Intel SGX and AMD SEV have been repeatedly broken by practical microarchitectural attacks (e.g., Foreshadow (Van Bulck et al., 2018), CacheOut (Van Schaik et al., 2021)) and fault-based exploits (e.g., Plundervolt (Murdock et al., 2020), SEV-Voltage Glitch (Buhren et al., 2021)). These attacks highlight a persistent mismatch between TEE threat models and real-world adversaries. Similarly, NVIDIA’s Confidential Computing (CC), e.g., H100 CC mode (Corporation, 2023), does not fully address well-resourced attackers. Recent work has reverse-engineered CC internals (Gu et al., 2025) and demonstrated practical breakages (Chuang et al., 2025) at low cost. Nvidia explicitly mentioned these attacks as out of scope: Advanced physical and side-channel attacks (Kiyan et al., 2024; Krachenfels et al., 2021; Mehta et al., 2025), limiting the robustness of CC-based solutions.
This lack of reliable visibility into utilization poses challenges for AI governance and safety. In the absence of trusted hardware telemetry, powerful GPUs can be silently repurposed to train or execute AI models for malicious purposes, including coordinated cyberattacks, automated vulnerability exploitation (Anthropic Team, 2025; Singh, 2025), advanced malware generation (Metta et al., 2024). At the same time, utilization monitoring must preserve user privacy and avoid leaking sensitive model architectures, training data, or application logic.
@@ 2. GPU in AI Workloads
This motivates telemetry-based utilization measurement, not as a substitute for attestation, but as a complementary signal for enforcement, anomaly detection, or fairness auditing. These mechanisms do not intend to provide cryptographic guarantees; instead, they expose observables such as compute rate, memory bandwidth, or solve-time distributions. Our measurements rely on no hardware trust anchors, but expose timing and memory patterns for future governance protocols.
@@ 3. Untrusted GPU Execution Settings
We consider a range of deployment cases in which external entities may wish to observe or estimate GPU usage for purposes such as telemetry logging, policy enforcement, or future integration into accountability protocols. The goal is to examine whether puzzle-based measurements (e.g., solve times, memory access penalties) provide a useful signal under minimal trust assumptions. The measurements we study are intended as heuristic indicators of usage, rather than full adversarial resistance or proofs. To this end, we define four representative settings that differ in which parts of the software/hardware stack can be assumed trustworthy.
@@ 4. Timing Measurements
This section introduces workload-based timing measurements that function as telemetry probes, providing soft-enforcement signals in untrusted environments without TPMs, TEEs, or trusted telemetry. These mechanisms aim to detect whether a worker (e.g., a GPU) has executed a specified unit of computation, when it occurred, and under what degree of contention. These measurements enable latency-based inference, which may inform governance-relevant mechanisms, such as job-scheduling integrity or co-location detection. Each workload probe reveals a different dimension of device behavior: tensor-core saturation (GEMM), parallel effort (PoW), or sequential pressure (VDF). A randomized challenge schedule enables these probes to be sampled over time, providing the challenger with temporal visibility into resource engagement.
@@ 4.1. Latency Model
Our approach leverages a measurement scheme rooted in memory-hard hash functions. We model PoW solve times under both single-thread and multi-thread (GPU) settings, showing the exponential-time behavior and memoryless property that enables a probabilistic mechanism. This method enforces a relationship between GPU utilization and the computational resources consumed during execution. By employing memory-hard primitives such as Argon2(Biryukov et al., 2016), the scheme ensures that a significant portion of the high-bandwidth memory (HBM) is actively engaged during the measurement process. A practical measurement procedure follows a pattern similar to use cases in cryptocurrencies: the challenger samples a strong random salt from a TRNG, composes a puzzle input that includes the salt and a trusted timestamp, and transmits it to the worker. The worker must perform a large-scale, parallel search or time-bound computation over the input (using thousands of GPU threads) and return a solution within the prescribed time window. The challenger validates the returned solution for correctness and checks that the observed response time lies within the expected interval. Deviations indicate possible under-utilization, outsourcing, or misbehavior. Although this mechanism introduces considerable power and thermal overhead, it provides a quantifiable effort metric of sustained memory utilization within the target GPU, making it well-suited for scenarios with the need for checking resource allocation. Shifting from PoW to GEMM can reduce power consumption and thermal overhead by aligning computation with native tensor-core execution, which is more efficient.
@@
The challenger samples two pseudorandom matrices derived from the salt and timestamp and instructs the worker to perform a sequence of dimension-controlled GEMM operations whose total FLOP count is tightly coupled to the requested utilization level. Because GEMM saturates both compute units and high-bandwidth memory, the resulting latency is highly reproducible on a local GPU but difficult to outsource to CPUs, small GPUs, or remote clouds. Moreover, the challenger can cheaply validate correctness by hashing the returned matrix or verifying random linear combinations of rows and columns. This GEMM-based PoW thus provides a practical, hardware-aligned mechanism for continuous utilization measurement that exploits the deterministic throughput characteristics of modern GPU matrix engines.
Sequential Delay. Shifting from PoW to Verifiable Delay Function (VDF) (Boneh et al., 2018; Wesolowski, 2020; Rivest et al., 1996) enables the challenger to capture not only parallel effort but also sequential compute pressure, offering finer-grained insight into whether the worker sustained uninterrupted execution over time. This capability is essential in governance scenarios where precomputation or parallel outsourcing must be disincentivized. To this end, we deploy a multi-instance VDF strategy that measures the effective sequential compute capacity of GPUs. Our design is optimized for GPU execution, minimizing kernel divergence and enhancing predictability. In particular, we adopt a construction that balances simplicity and efficiency (Wesolowski, 2020), which is well-suited for continuous utilization monitoring in a multi-instance setting across thousands of GPU cores. Each VDF instance is solved independently and sequentially, ensuring that parallelism provides throughput gains only within controlled architectural limits. At regular intervals, the challenger generates randomized VDF puzzles and transmits each input, including a securely generated salt and timestamp, to the worker. The GPU then performs multi-instance VDF evaluations locally by executing sequential squaring operations over large modular arithmetic groups. Upon completion, the worker returns both the computed outputs and associated proof elements to the challenger. The challenger then checks both the correctness of the computation and the timing constraints associated with the response. To mitigate the risk of outsourcing to alternate platforms such as CPU clusters or ASICs, one tunes instance difficulty, modulus size, and iteration depth to favor GPU architectures.
@@ 4.2. Continuous Compute-based Measurement
The challenger continuously sends the worker randomized, time-bound puzzles, enforces an expected utilization level by adjusting difficulty probabilistically, and resists trivial outsourcing by building puzzles that exploit GPU architectural advantages (massive parallelism, tensor cores, and high memory bandwidth) via memory-hard functions, GEMMs, and multi-instance VDFs. Verification checks both solution validity and timing to detect overuse or policy violations.
@@ 4.3. Random-Timed Challenge-Response
In contrast to the continuous approach, which incurs inherent unavoidable power and computation overhead, we propose a random-timed Challenge-Response mechanism. For this purpose, we focus on the architectural features that are particularly ly belonged to GPUs. Specifically we build upon our approach based on GPU’s high-bandwidth and large memory. Furthermore, we emphasis the heavy memory transitions in today’s LLMs on cutting edge GPUs and select the memory utilization for measurement.
Figure 2 shows a high-level flow of the proposed challenge–response mechanism, which explicitly exploits the properties of the GPU’s high-bandwidth memory (HBM). In this scheme, a pre-determined challenge data block is stored inside the GPU’s VRAM before the measurement process begins. This part, highlighted in red in Figure 2 and referred to as the Init. Mem., serves as the trusted memory base for all subsequent computations. Any violation of this storage requirement (e.g., moving challenge data off HBM) invalidates the timing assumptions. Our detection method relies on the observation that the computation delay of a given hash function operating on data locally stored in HBM is statistically lower than when the same data must traverse off-chip over PCIe. This latency gap forms the basis for identifying compliant versus non-compliant behavior.
@@ 4.4. Comparison of the proposed methods
The CRP method leverages the architectural properties of the GPU’s High Bandwidth Memory (HBM) rather than relying solely on computation-related assumptions. By verifying whether challenge data resides locally in HBM, it detects off-chip data accesses with minimal delay, achieving low detection time and negligible continuous power overhead. While it primarily offers architectural rather than compute-based enforcement, it provides strong resistance to outsourcing and enables near-instant anomaly detection via timing deviations between on-chip and PCIe-bound operations.
@@ 6.4. Overhead Analysis
This follows directly from the proposed design where the CHAL region must occupy a large, contiguous portion of the GPU’s HBM so that residency violations can be detected with high sensitivity. Consequently, the memory footprint of the VRAM Residency test is large by design, even though it performs no additional allocations beyond the CHAL itself.
@@ 7. Conclusion and Discussion
While our fingerprinting-based masking mechanism was evaluated in a non-continuous setting, its integration with randomized probes offers a promising direction for future continuous telemetry mechanisms. Periodic sampling with ephemeral nonces and small Argon2 probes could enforce ongoing residency checks with low overhead, subject to constraints imposed by co-residency, scheduling, and bandwidth fairness. Designing such extensions without disrupting legitimate workloads remains an open challenge. While Section 6 shows separable timing distributions, defining a hardware-specific formal thresholds left for the future work. As such, FP/FN rates are not quantified, though they can be estimated empirically under deployment-specific assumptions. A detection-theoretic framework remains important future work.